OpenStudy (anonymous):

Solve the following equation. Please show work. -10-3/7x=5

OpenStudy (anonymous):

first simplify -10-3=-13..so -13/7x=5 multiply each side by 7x -13 = 35x divide by 35 on each side x=-13/35